OpenStudy (volleyballlover55):

A square rug covers 79 square feet of floor. What is the approximate length of one side of the rug? (Approximate to the nearest hundredth foot.) 8.86 feet 8.87 feet 8.88 feet 8.89 feet

OpenStudy (volleyballlover55):

@Astrophysics

jhonyy9 (jhonyy9):

how you get area of a square ?

OpenStudy (calculusxy):

 A square has all of its sides equal to each other (in other words known as congruent). Thus, if one side, for example, is equal to 8in., then all of the other sides are 8in. So to find the area of the square, you multiply length x width (in this case 8in. x 8in). What you really have, in simpler terms, \(A = s^2\).

OpenStudy (volleyballlover55):

The area of a square can be found by multiplying the base times itself. This is similar to the area of a rectangle but the base is the same length as the height.

OpenStudy (calculusxy):

`The area of a square can be found by multiplying the base times itself` This part is very important. From this you derive the formula \(Area = s^2\)

OpenStudy (volleyballlover55):

ok

OpenStudy (calculusxy):

What is the area of the square rug?

OpenStudy (volleyballlover55):

umm.. 16?

OpenStudy (volleyballlover55):

@calculusxy

jhonyy9 (jhonyy9):

than you know now that area of a square is side time side and you know that area what is given is equal 79 A = 79 = s^2 s = ?

OpenStudy (volleyballlover55):

8?

OpenStudy (volleyballlover55):

or 64?

OpenStudy (jchick):

It is as easy as saying what times what equals 79 or close to it. Check your answers by multiplying by themselves.

OpenStudy (jchick):

8.86 * 8.86 is 78.4996

OpenStudy (jchick):

8.87 * 8.87 is 78.6769

OpenStudy (jchick):

8.88 * 8.88 is 78.8544

OpenStudy (jchick):

8.89 * 8.89 is 79.0321

OpenStudy (jchick):

So you can see your answer is D.

OpenStudy (volleyballlover55):

yay! okay thx :)

OpenStudy (volleyballlover55):

can you help me with some more questions please? :)

OpenStudy (jchick):

Because like they have said above to find the area of a square is it A = s^2

OpenStudy (jchick):

Sure!

OpenStudy (volleyballlover55):

ok ill tag you :)

OpenStudy (jchick):

Ok